if a car rented for $23 per day for 3 days, and needed to pay $2.5 for each mile if the total cost was $150 for renting the car how many miles can be driven for 3 days?
Answer by Maths68(1474)

Assume, miles can be driven for 3 days = x
Given
Car rent/day = 23
Rent of three day = 3 * 23 = 69
For a mile we needed to pay = 2.5
Total rent for driving x miles = 2.5(x)
Total amount paid = 150
Total rent for driving x miles + Rent of three day = Total amount paid
2.5(x)+69=150
2.5(x)=150-69
2.5(x)=81
Divide by 2.5 both sides of the equation
2.5(x)/2.5=81/2.5
x=32.4
32.4 miles can be driven for 3 days.